# Hierarchical Relation-augmented Representation Generalization for Few-shot Action Recognition

### v1

@@ 4.4. Diagnostic Experiment
Table 6. Generalization performance of our method with different temporal alignment metrics on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the $5$-way $1$-shot setting (§4.6).
@@
4.5. Quality Analysis
Similarity visualization. To qualitatively demonstrate the effectiveness of hierarchical relation modeling in our HR2G-shot, we visualize the predicted similarities between query and support prototypes with different approaches (i.e., CLIP-FSAR (Wang et al., 2024b) and ours) for one task of HMDB51 in Fig. 5. As seen, our HR2G-shot can make more accurate decisions for similar classes in each task compared with CLIP-FSAR. Specifically, for the first query sample in Fig. 5, the incorrect decision obtained by CLIP-FSAR can be rectified by jointly modeling three types of relations. These results further demonstrate that our method can capture task-specific temporal cues via transferring knowledge at both video-level and task-level.
The visualization of Feature Distribution. Fig. 6 visualizes learned features of baseline and our algorithm HR2G-shot with ISC and IKT via the t-SNE tool. We clearly observe that after conducting inter-video and inter-task interactions, learned video features become more compact (intra-class) and better separated (inter-class). As shown in Fig. 6(b), the three classes “Smoke”, “Sit”, and “Pick” become clearly distinguishable from each other after conducting hierarchical relation modeling. The above phenomena suggest that we can gain task-specific temporal features by transferring knowledge at both video-level and task-level, ensuring better feature discrimination. See more examples in Appendix.
4.6. Generalization Study
To demonstrate that HR2G-shot generalizes well to different alignment metrics, Table 6 report comparison results on HMDB51 (Kuehne et al., 2011) and SSv2-small (Goyal et al., 2017) with regard to different temporal metrics (cf. Eq. 7). Different temporal metrics are adopted, including OTAM (Cao et al., 2020), TRX (Perrett et al., 2021) and Bi-MHM (Wang et al., 2022). As seen, our HR2G-shot consistently achieves better performance compared with CLIP-FSAR (Wang et al., 2024b), regardless of the temporal alignment metric employed. This shows that our method can adapt to any temporal alignment metric.
5. Conclusion
In this work, we present HR2G-shot, a novel Hierarchical Relation-augmented Representation Generalization framework for FSAR, which captures task-specific temporal cues from multiple perspectives (i.e., inter-frame, inter-video, and inter-task). Rather than only conducting inter-frame temporal interactions, we further i) perform fine-grained cross-video temporal interaction for exploring inter-video relationships, and ii) aggregate useful temporal knowledge from previous tasks for learning inter-task relationships. Our framework is hierarchical, elegant, and gains outstanding performance on five standard FSAR datasets, i.e., SSv2-full, SSv2-small, HMDB51, UCF101, and Kinetics. Our work offers a fresh perspective to current FSAR solutions by hierarchically exploring multi-level relations, and we wish it to inspire future research in this field.

### v3

@@ 4.6.2 Results with Various Ways and Shots
Table 12: Generalization performance of our method with different temporal alignment metrics on SSv2-small [71] and HMDB51 [64] under the 55-way 11-shot setting (§4.6.3).
@@
Table 13: Comparison with SOTAs using the ResNet-based CLIP backbone under the 5-way 1-shot setting (§4.6.1).
  Method SSv2-small UCF101 HMDB51
CLIP-FSAR [61] $52.0$ $91.3$ $72.5$
MVP-shot [68] $51.2$ $92.2$ $72.5$
HR2G-shot (ours) $\mathbf{53.1}$ $\mathbf{93.7}$ $\mathbf{74.5}$
4.6.3 Results with Various Metrics
To demonstrate that HR2G-shot generalizes well to different alignment metrics, Table 12 report comparison results on HMDB51 [64] and SSv2-small [71] with regard to different temporal metrics (cf. Eq. 7). Different temporal metrics are adopted, including OTAM [10], TRX [19] and Bi-MHM [13]. The results clearly indicate that the benefits of our HR2G-shot are not limited to a specific alignment strategy. Our method consistently outperforms CLIP-FSAR by significant margins across all settings. Notably, while the baseline performance fluctuates depending on the metric, our method maintains high accuracy (e.g., $\sim$$60\%$ on SSv2-small and $\sim$$85\%$ on HMDB51). This demonstrates that the hierarchical relation modeling in our framework learns intrinsically better video representations, thereby unlocking the potential of downstream alignment metrics.
5 Conclusion and Discussion
Conclusion. In this work, we present HR2G-shot, a novel Hierarchical Relation-augmented Representation Generalization framework for FSAR, which captures task-specific temporal cues from multiple perspectives (i.e., inter-frame, inter-video, and inter-task). Rather than only conducting inter-frame temporal interactions, we further i) perform fine-grained cross-video temporal interaction for exploring inter-video relationships, and ii) aggregate useful temporal knowledge from previous tasks for learning inter-task relationships. Our framework is hierarchical, elegant, and gains outstanding performance on five standard FSAR datasets, i.e., SSv2-full, SSv2-small, HMDB51, UCF101, and Kinetics. Our work offers a fresh perspective to current FSAR solutions by hierarchically exploring multi-level relations, and we wish it to inspire future research in this field.
Discussion. One limitation of our algorithm is that it retrieves and aggregates temporal knowledge from historical tasks, yet ignores spatial information in actions. In the future, we plan to upgrade the knowledge bank to jointly encode and transfer task-specific spatio-temporal patterns. Additionally, while our work primarily focuses on few-shot action recognition tasks, the proposed hierarchical relational paradigm possesses high versatility. In the future, we aim to extend this framework to a broader range of applications, including zero-shot action recognition and other related tasks.